# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 5. Sylow Invariants
The order of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ is $2q^{m(m-1)}(q^{m}-1)\prod_{j=1}^{m-1}(q^{2j}-1)$, see [23, page 141]. Therefore a Sylow $p$-subgroup has order $q^{m(m-1)}$.
The field of fractions for the Sylow invariants was calculated in [14] and the ring of invariants for the $m=2$ case was calculated in [15].
@@
$y_{1}\cdot h=y_{1}+\sum_{i=2}^{m}(b_{i}y_{i}+a_{i}x_{i}-a_{i}b_{i}x_{1}),$
$x_{1}\cdot h=x_{1}$, and for $j>1$, $y_{j}\cdot h=y_{j}-a_{j}x_{1}$, $x_{j}\cdot h=x_{j}-b_{j}x_{1}$ where $a_{i},b_{i}\in\mathbb{F}_{q}$. Hence the order of $H$ is $q^{2m-2}$. Note that $H$ is a normal subgroup of $P_{m}$. Identify $P_{m-1}$ with pointwise stabiliser of $\{y_{1},x_{1}\}$ in $P_{m}$. Then $P_{m}=P_{m-1}H$. (Note that $(m-1)(m-2)+(2m-2)=m(m-1)$.) Therefore $S_{m}^{P_{m}}=(S_{m}^{H})^{P_{m-1}}$. We will calculate $S_{m}^{P_{m}}$ by first computing the Hook invariants.
@@ 5.1. Computing the Hook Invariants
$Q:=\mathbb{F}_{q}[y_{2},\ldots,y_{m},x_{m},\ldots,x_{1}]\subset S_{m}.$
For $i>1$, define $Y_{i}:=y_{i}^{q}-y_{i}x_{1}^{q-1}$ and $X_{i}=x_{i}^{q}-x_{i}x_{1}^{q-1}$. It is easy to see that
$Q^{H}:=\mathbb{F}_{q}[x_{1}][Y_{i},X_{i}\mid i=2,\ldots,m].$
Since $\xi_{0}$ has degree $1$ in $y_{1}$, we have $S_{m}^{H}[x_{1}^{-1}]=Q^{H}[\xi_{0},x_{1}^{-1}]$ (see [3, Theorem 2.4]).
A straight-forward calculation gives
(1) $\xi_{0}^{q}-\xi_{1}x_{1}^{q-1}+\xi_{0}x_{1}^{2q-2}=\sum_{i=2}^{m}X_{i}Y_{i}=\psi_{1}(\xi_{0}).$
For $j>0$, further calculation, or the application of Steenrod operations, gives
(2) $\xi_{j}^{q}-\xi_{j+1}x_{1}^{q-1}=\xi_{1}x_{1}^{q^{j+1}-1}-2\xi_{0}x_{1}^{q^{j+1}+q-2}+\psi_{1}(\xi_{j}).$
Observe that the orbit product of $y_{1}$ over $H$ coincides with the orbit product of $y_{1}$ over $P_{m}$.
@@
$x_{1}N(y_{1})-\xi_{n-2}\in Q^{H}[\xi_{i}\mid i=0,\ldots,n-3]$
@@
$A:=Q^{H}[N(y_{1}),\xi_{j}\mid j=0,\ldots,n-3]$
and $\mathcal{W}:=\{X_{i},Y_{i}\mid i=2,\ldots,m\}$.
@@
$\mathcal{W}\cup\{N(y_{1}),x_{1},\xi_{j}\mid j=0,\ldots,n-3\}$
is a Khovanskii basis for $A$ using the lexicographic order. Furthermore, $A$ is a complete intersection with relations given by the subductions of the tête-a-têtes $\xi_{n-3}^{q}-x_{1}^{q}N(y_{1})$ and $\xi_{j}^{q}-x_{1}^{q-1}\xi_{j+1}$ for $j=0,\ldots,n-4$.
@@
The tête-a-têtes $\xi_{j}^{q}-\xi_{j+1}x_{1}^{q-1}$ for $j=1,\ldots,n-4$ subduct to zero using Equations 1 and 2. The subduction of the tête-a-tête $x_{1}^{q}N(y_{1})-\xi_{n-3}^{q}$ is constructed using Equation 2 with $j=n-3$ and Lemma 5.2 to substitute for $x_{1}^{q-1}\xi_{n-2}$. ∎
@@
Since $Q^{H}[N(y_{1})]$ is generated by a homogeneous system of parameters and $S_{m}^{H}[x_{1}^{-1}]=Q^{H}[\xi_{0},x_{1}^{-1}]$, it is sufficient to prove that the ring $A$ is integrally closed in its field of fractions. We do this by proving $x_{1}$ is prime in $A$ which is sufficient by [19, Theorem 20.2].
@@
It follows from Lemma 5.3 that $A$ is Cohen-Macaulay and has rank $q^{n-2}$ over $Q^{H}[N(y_{1})]$ with a basis given by the monomial factors of $(\xi_{0}\cdots\xi_{n-3})^{q-1}$. We will show that $x_{1}$ is prime in $A$ by proving that $A/x_{1}A$ is an integral domain.
Recall that $\psi_{1}:S_{m}\to S_{m}$ is the algebra homomorphism determined by $\psi_{1}(w)=w^{q}-wx_{1}^{q-1}$ for $w$ homogeneous of degree one. Clearly $x_{1}S_{m}\subset\ker(\psi_{1})$. To see that $x_{1}S_{m}=\ker(\psi_{1})$, identify $S_{m}/x_{1}S_{m}$ with $\mathbb{F}_{q}[y_{1},\ldots,x_{2}]$ and observe that $\psi_{1}$ maps this ring injectively to $\mathbb{F}_{q}[y_{1}^{q}-y_{1}x_{1}^{q-1},Y_{2},\ldots,X_{2}]$.
@@
Let $\overline{\psi_{1}}$ denote the algebra map from $A/x_{1}A$ to $S_{m}$ induced by $\psi_{1}$. If $\overline{\psi_{1}}$ is injective, then $A/x_{1}A$ is an integral domain, proving $x_{1}A$ is a prime ideal and $A$ is normal.
@@
Define $\overline{Q}=\mathbb{F}_{q}[N(y_{1}),Y_{2},\ldots,X_{2}]$, i.e., the subalgebra generated by $\{N(y_{1})\}\cup\mathcal{W}$. Let $\mathcal{B}_{H}$ denote the set of monomial factors of $(\xi_{0}\cdots\xi_{n-3})^{q-1}$. Using the $Q^{H}$-module structure of $A$, we see that $A/x_{1}A$ is a free $\overline{Q}$-module with a basis $\mathcal{B}_{H}$. By identifying $\overline{Q}$ with the submodule of $A/x_{1}A$ generated by $1$ we see that $A/x_{1}A$ is a complete intersection with relations $\xi_{i}^{q}=\psi_{1}(\xi_{i})$ for $i=0,\ldots,n-3$. Observe that for $w\in S_{m}$ with $\deg(w)=1$, $\psi_{1}\circ\psi_{1}(w)=\psi_{1}(w)^{q}$. Therefore $\psi_{1}(A)$ is a module over $\mathbb{F}_{q}[\psi_{1}(N(y_{1})),Y_{2}^{q},\dots,X_{2}^{q}]$ with module generators $\{\psi_{1}(\beta)\mid\beta\in\mathcal{B}_{H}\}$. To prove that $\overline{\psi_{1}}$ is injective, it is sufficient to show that $\overline{\psi_{1}}(A/x_{1}A)=\psi_{1}(A)$ is a free module over $\mathbb{F}_{q}[\psi_{1}(N(y_{1})),Y_{2}^{q},\dots,X_{2}^{q}]$ with module generators $\{\psi_{1}(\beta)\mid\beta\in\mathcal{B}_{H}\}$. Note that for $\beta\in\mathcal{B}_{H}$, we have $\psi_{1}(\beta)\in\mathbb{F}_{q}[Y_{2}^{q},\dots,X_{2}^{q}]$. Therefore, it is sufficient to show that the $\mathbb{F}_{q}[Y_{2}^{q},\dots,X_{2}^{q}]$-module generated by $\{\psi_{1}(\beta)\mid\beta\in\mathcal{B}_{H}\}$ is free. After relabelling and a reindexing which takes $n-2$ to $n$, this is equivalent to showing that the $\mathbb{F}_{q}[y_{1}^{q},\dots,x_{1}^{q}]$-module generated by $\{\xi_{0}^{e_{0}}\cdots\xi_{n-1}^{e_{n-1}}\mid e_{i}<q\}\subset S_{m}$ is free of rank $q^{n}$.
Let $G$ denote a finite orthogonal group with defining representation $V$. We know that the invariant field $\mathbb{F}_{q}(V)^{G}$ is generated by $\{\xi_{0},\ldots,\xi_{n-1}\}$. Since the extension $\mathbb{F}_{q}(\xi_{0},\ldots,\xi_{n-1})=\mathbb{F}_{q}(V)^{G}\subset\mathbb{F}_{q}(V)$ is a Galois extension, for any intermediate extension $\mathbb{F}_{q}(V)^{G}\subset E\subset\mathbb{F}_{q}(V)$, we have $E=\mathbb{F}_{q}(V)^{L}$ for some subgroup $L\leq G$. Consider $E=\mathbb{F}_{q}(V)^{G}[y_{1}^{q},...,x_{1}^{q}]$. Since the pointwise stabiliser of $\{y_{1}^{q},...,x_{1}^{q}\}$ in ${{\text{GL}_{{n}}(\mathbb{F}_{q})}}$ is $\{1\}$, we have $\mathbb{F}_{q}(V)^{G}[y_{1}^{q},...,x_{1}^{q}]=\mathbb{F}_{q}(V)$. On the other hand, $\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})\subset\mathbb{F}_{q}(V)$ is a field extension of degree $q^{n}$. Construct this extension by iteratively adjoining the $\xi_{i}$. Since $\xi_{i}$ is a root of $t^{q}-\xi_{i}^{q}\in\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})[t]$, at each step the extension has degree at most $q$. To complete the construction in $n$ steps, we need each step to be an extension of degree $q$. We can adjoin the $\xi_{i}$ in any order. Therefore $t^{q}-\xi_{i}^{q}$ is irreducible over $\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})[\xi_{j}\mid j\not=i]$. Hence $\{\xi_{0}^{e_{0}}\cdots\xi_{n-1}^{e_{n-1}}\mid e_{i}<q\}$ is a basis for $\mathbb{F}_{q}(V)$ as a vector space over $\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})$. This means that $\{\xi_{0}^{e_{0}}\cdots\xi_{n-1}^{e_{n-1}}\mid e_{i}<q\}$ is linearly independent over $\mathbb{F}_{q}(y_{1}^{q},...,x_{1}^{q})$ and generates a free $\mathbb{F}_{q}[y_{1}^{q},...,x_{1}^{q}]$-module of rank $q^{n}$, completing the proof. ∎
The rest of Section 5.1 is devoted to the proof of Lemma 5.2. For a non-zero element $h\in S=S_{m}$ we denote by $\nu_{1}(h)$ the greatest integer $n$ such that $h\in(x_{1})^{n}S$, i.e., $x_{1}^{\nu_{1}(h)}$ divides $h$ in $S$ but $x_{1}^{\nu_{1}(h)+1}$ does not. Let $d$ be a non-negative integer. We write $||d||_{q}=||d||$ to denote the digit sum in base $q$, i.e., if $d=\sum_{i=0}^{r}d_{i}q^{i}$ with $0\leq d_{i}<q$ for all $i$ then $||d||_{q}=\sum_{i=0}^{r}d_{i}$. We will say that a polynomial $f=f(y_{1})=\sum_{k}c_{k}y_{1}^{k}\in S$ with each $c_{k}\in Q$ is compliant (with respect to $x_{1}$) if $\nu_{1}(c_{k})\geq||k||-1$ for all $k\geq 1$. We say that $f$ is strongly compliant if $\nu_{1}(c_{k})\geq||k||$ for all $k\geq 1$. Observe that $\xi_{i}$ is strongly compliant for all $i$ and that the set of strongly compliant polynomials is closed under addition. Furthermore, since $||k+\ell||\leq||k||+||\ell||$, the set of strongly compliant polynomials is closed under multiplication.
@@
If $f\in S^{H}$ is strongly compliant and $\deg_{y_{1}}(f)<q^{n-2}$, then $f\in Q^{H}[\xi_{i}\mid i=0,\ldots,n-3]$.
@@
The proof is by induction on $d:=\deg_{y_{1}}(f)$. For $d=0$, $f=c_{0}\in Q^{H}$. Suppose the result is true for all strongly compliant invariants with $0<d<\ell<q^{n-2}$.
Consider $f=\sum_{k=0}^{\ell}c_{k}y_{1}^{k}\in S^{H}$ with $c_{k}\in Q$ and $c_{\ell}\not=0$. Since $y_{1}-y_{1}\cdot g\in Q$ for all $g\in H$, it follows that $c_{\ell}\in Q^{H}$. Write $\ell=\sum_{i=0}^{n-3}\ell_{i}q^{i}$ with $0\leq\ell_{i}<q$. Define $\beta:=\prod_{j=0}^{n-3}\xi_{j}^{\ell_{j}}$. Then $\operatorname{LT}(\beta)=y_{1}^{\ell}x_{1}^{||\ell||}.$ Since $f$ is strongly compliant, $x_{1}^{||\ell||}$ divides $c_{\ell}$. Define $h:=f-c_{\ell}x_{1}^{-||\ell||}\beta$. Observe that $h\in S^{H}$. Since $\beta$ is strongly compliant, $h$ is strongly compliant with $\deg_{y_{1}}(h)<\ell$. Therefore, using the induction hypothesis, $h\in Q^{H}[\xi_{i}\mid i=0,\ldots,n-3]$. Thus $f=h+c_{\ell}x_{1}^{-||\ell||}\beta\in Q^{H}[\xi_{i}\mid i=0,\ldots,n-3]$. ∎
We will prove by induction on $m$ that $N(y_{1})$ is compliant. As a consequence, $x_{1}N(y_{1})-\xi_{n-2}$ is strongly compliant. Lemma 5.2 then follows from Lemma 5.5.
We first observe that $N(y_{1})$ is compliant for $m=1$. For ${{\text{O}^{+}_{{2}}(\mathbb{F}_{q})}}$, we have $N(y_{1})=y_{1}$, which is trivially compliant.
@@
Suppose that $q-1$ divides $k$, then $q-1$ divides $||k||_{q}$.
@@
Write $k=\sum_{s}k_{s}q^{s}$ where $10\leq k_{s}\leq q-$ 1 for all $s$. Since $q^{s}\equiv 1\pmod{(q-1)}$ for all $s$ we have
$0\equiv k=\sum_{s}k_{s}q^{s}\equiv\sum_{s}k_{s}=||k||_{q}\pmod{(q-1)}$
@@
Define $T:=S[t,u]$ and extend the definition of compliance to $T$. Given $f(y_{1},t)\in T$, we define
$\mathcal{T}_{(y_{1},u,t,x_{1})}(f):=\prod_{a\in\mathbb{F}_{q}}f(y_{1}+au,t-ax_{1})\in T\ .$
@@
If $f$ is compliant then $\mathcal{T}_{(y_{1},u,t,x_{1})}(f)$ is compliant.

### v3

@@ 5. Sylow Invariants
The order of $\text{O}^{+}_{{2m}}(\mathbb{F}_{q})$}} is $2q^{m(m-1)}(q^{m}-1)\prod_{j=1}^{m-1}(q^{2j}-1)$, see [27, page 141]. Therefore a Sylow $p$-subgroup has order $q^{m(m-1)}$. The field of fractions for the Sylow invariants was calculated in [14] and the ring of invariants for the $m=2$ case was calculated in [15].
@@
$y_{m}\cdot h=y_{m}+\sum_{i=1}^{m-1}(b_{i}y_{i}+a_{i}x_{i}-a_{i}b_{i}x_{m}),$
$x_{m}\cdot h=x_{m}$, and for $j<m$, $y_{j}\cdot h=y_{j}-a_{j}x_{m}$, $x_{j}\cdot h=x_{j}-b_{j}x_{m}$ where $a_{i},b_{i}\in\mathbb{F}_{q}$. Hence the order of $H$ is $q^{2m-2}$. Note that $H$ is a normal subgroup of $P_{m}$. Identify $P_{m-1}$ with pointwise stabiliser of $\{y_{m},x_{m}\}$ in $P_{m}$. Then $P_{m}=P_{m-1}H$. (Note that (m1)(m2)+(2m2)=m(m1)(m-1)(m-2)+(2m-2)=m(m-1).) Therefore $S_{m}^{P_{m}}=(S_{m}^{H})^{P_{m-1}}$. We will calculate $S_{m}^{P_{m}}$ by first computing the Hook invariants. Since the orbit product of $y_{m}$ over $H$ coincides with the orbit product of $y_{m}$ over $P_{m}$, we have $N(y_{m})=N^{H}(y_{m})$ where $N^{H}(a)$ is the orbit product of $a$ over $H$.
@@ 5.1. Computing the Hook Invariants
$Q:=\mathbb{F}_{q}[y_{m-1},\ldots,y_{1},x_{1},\ldots,x_{m}]\subset S_{m}.$
For $i<m$, define $Y_{i}:=y_{i}^{q}-y_{i}x_{m}^{q-1}$ and $X_{i}=x_{i}^{q}-x_{i}x_{m}^{q-1}$. It is easy to see that
$Q^{H}:=\mathbb{F}_{q}[x_{m}][Y_{i},X_{i}\mid i=1,\ldots,m-1].$
Since $\xi_{1}$ has degree $1$ in $y_{m}$, we have $S_{m}^{H}[x_{m}^{-1}]=Q^{H}[\xi_{1},x_{m}^{-1}]$ (see [3, Theorem 2.4]).
Using the lex order $\operatorname{LT}(\xi_{j}^{q})=x_{m}^{q}y_{m}^{q^{j}}=\operatorname{LT}(\xi_{j+1}x_{m}^{q-1})$. To subduct the tête-a-tête $\xi_{j}^{q}-\xi_{j+1}x_{m}^{q-1}$, we first eliminate $y_{m}$. For $j=1$, we get
(4) $\xi_{1}^{q}-\xi_{2}x_{m}^{q-1}+\xi_{1}x_{m}^{2q-2}=\psi_{[m,1]}(\xi_{1})\in Q^{H}$
(compare with Equation 1). For $j>1$, we have
$\xi_{j}^{q}-\xi_{j+1}x_{m}^{q-1}-\xi_{1}^{q}x_{m}^{q^{j}-q}+\xi_{1}x_{m}^{q^{j}+q-2}\in Q^{H}.$
Since $Q^{H}$ is a polynomial algebra whose generators have relatively prime lead terms, the subduction algorithm can be used to write an element of $Q^{H}$ as a polynomial in the generators. Therefore the tête-a-tête $\xi_{j}^{q}-\xi_{j+1}x_{m}^{q-1}$ subducts to zero in $Q^{H}[\xi_{1},\ldots,\xi_{j+1}]$. For $j=1$, the subduction is given by Equation 4. For $j>1$, the subduction is given by the following lemma.
@@
For $j>1$, we have
$\xi_{j}^{q}-\xi_{j+1}x_{m}^{q-1}=\xi_{1}^{q}x_{m}^{q^{j}-q}-\xi_{1}x_{m}^{q^{j}+q-2}+\sum_{i=1}^{j}\psi_{[m,1]}(\xi_{i})x_{m}^{q^{j}-q^{i}}.$
Proof.
The proof is by induction on $j$. Using Equation 2 and the fact that $\psi$ followed by evaluating $t$ at $x_{m}$ gives $\psi_{[m,1]}$, we have
$\xi_{2}^{q}-\xi_{3}x_{m}^{q-1}$ $=$ $2\xi_{1}^{q}x_{m}^{q^{2}-q}-\xi_{2}x_{m}^{q^{2}-1}+\psi_{[m,1]}(\xi_{2})$
$=$ $\psi_{[m,1]}(\xi_{2})+\xi_{1}^{q}x_{m}^{q^{2}-q}+x_{m}^{q^{2}-q}(\xi_{1}^{q}-\xi_{2}x_{m}^{q-1}).$
Using Equation 4 to substitute for $\xi_{1}^{q}-\xi_{2}x_{m}^{q-1}$ gives
$\xi_{2}^{q}-\xi_{3}x_{m}^{q-1}$ $=$ $\psi_{[m,1]}(\xi_{2})+\xi_{1}^{q}x_{m}^{q^{2}-q}+x_{m}^{q^{2}-q}(\psi_{[m,1]}(\xi_{1})-\xi_{1}x_{m}^{2q-2})$
$=$ $\xi_{1}^{q}x_{m}^{q^{2}-q}-\xi_{1}x_{m}^{q^{2}+q-2}+\psi_{[m,1]}(\xi_{2})+x_{m}^{q^{2}-q}\psi_{[m,1]}(\xi_{1}),$
proving the result for $j=2$.
For $j>2$, Equation 3 gives
$\xi_{j}^{q}-\xi_{j+1}x_{m}^{q-1}$ $=$ $\xi_{j-1}^{q}x_{m}^{q^{j-1}(q-1)}-\xi_{j}x_{m}^{(q^{j-1}+1)(q-1)}+\psi_{[m,1]}(\xi_{j})$
$=$ $\psi_{[m,1]}(\xi_{j})+x_{m}^{q^{j-1}(q-1)}(\xi_{j-1}^{q}-\xi_{j}x_{m}^{q-1}).$
Using the induction hypothesis
$x_{m}^{q^{j-1}(q-1)}(\xi_{j-1}^{q}-\xi_{j}x_{m}^{q-1})$ $=$ $x_{m}^{q^{j-1}(q-1)}(\xi_{1}^{q}x_{m}^{q^{j-1}-q}-\xi_{1}x_{m}^{q^{j-1}+q-2}$
$+\sum_{i=1}^{j-1}\psi_{[m,1]}(\xi_{i})x_{m}^{q^{j-1}-q^{i}})$
$=$ $\xi_{1}^{q}x_{m}^{q^{j}-q}-\xi_{1}x_{m}^{q^{j}+q-2}+\sum_{i=1}^{j-1}\psi_{[m,1]}(\xi_{i})x_{m}^{q^{j}-q^{i}}.$
Therefore
$\xi_{j}^{q}-\xi_{j+1}x_{m}^{q-1}=\xi_{1}^{q}x_{m}^{q^{j}-1}-\xi_{1}x_{m}^{q^{j}+q-2}+\sum_{i=1}^{j}\psi_{[m,1]}(\xi_{i})x_{m}^{q^{j}-q^{i}},$
as required. ∎
Lemma 5.3.
$x_{m}N(y_{m})-\xi_{n-1}\in Q^{H}[\xi_{1},\ldots,\xi_{n-2}]$.
@@
$A:=Q^{H}[\xi_{1},\ldots,\xi_{n-2}][N(y_{m})]$
and $\mathcal{W}:=\{X_{i},Y_{i}\mid i=1,\ldots,m-1\}$.
@@
$\mathcal{W}\cup\{N(y_{m}),x_{m},\xi_{1},\dots,\xi_{n-2}\}$
is a Khovanskii basis for $A$ using the lexicographic order. Furthermore, $A$ is a complete intersection with relations given by the subductions of the tête-a-têtes $\xi_{n-2}^{q}-x_{m}^{q}N(y_{m})$ and $\xi_{j}^{q}-x_{m}^{q-1}\xi_{j+1}$ for $j=1,\ldots,n-3$.
@@
The tête-a-têtes $\xi_{j}^{q}-x_{m}^{q-1}\xi_{j+1}$ for $j=1,\ldots,n-3$ subduct to zero using Equation 4 and Lemma 5.2. The subduction of the tête-a-tête $x_{m}^{q}N(y_{m})-\xi_{n-2}^{q}$ is constructed using Lemma 5.3 and Lemma 5.2 with j=n2j=n-2. Using Lemma 5.3, define $F:=x_{m}N(y_{m})-\xi_{n-1}\in Q^{H}[\xi_{1},\ldots,\xi_{n-2}]$ so
$\xi_{n-2}^{q}-x_{m}^{q}N(y_{m})=\xi_{n-2}^{q}-x_{m}^{q-1}\xi_{n-1}-Fx_{m}^{q-1}.$
Using Lemma 5.2 with j=n2j=n-2 gives
$\xi_{n-2}^{q}-x_{m}^{q}N(y_{m})=-Fx_{m}^{q-1}+\xi_{1}^{q}x_{m}^{q^{n-2}-q}-\xi_{1}x_{m}^{q^{n-2}+q-2}+\sum_{i=1}^{n-2}\psi_{[m,1]}(\xi_{i})x_{m}^{q^{n-2}-q^{i}}$
with $\psi_{[m,1]}(\xi_{i})\in Q^{H}$. Subduction can be used to write $F$ in terms of the generators of $Q^{H}[\xi_{1},\ldots,\xi_{n-2}]$, completing the calculation.
Since the non-trivial tête-a-têtes subduct to zero, we have a Khovanskii basis. When we have a Khovanskii basis for an algebra, the ideal of relations is generated by the subductions of the non-trivial tête-a-têtes. Therefore $A$ is a complete intersection. ∎
@@
Since $Q^{H}[N(y_{m})]$ is generated by a homogeneous system of parameters and $S_{m}^{H}[x_{m}^{-1}]=Q^{H}[\xi_{1},x_{m}^{-1}]$, it is sufficient to prove that the ring $A$ is integrally closed in its field of fractions. We do this by proving $x_{m}$ is prime in $A$ which is sufficient by [21, Theorem 20.2].
@@
It follows from Lemma 5.4 that $A$ is Cohen-Macaulay and has rank $q^{n-2}$ over $Q^{H}[N(y_{m})]$ with a basis given by the monomial factors of $(\xi_{1}\cdots\xi_{n-2})^{q-1}$. We will show that $x_{m}$ is prime in $A$ by proving that $A/x_{m}A$ is an integral domain.
@@
Recall that $\psi_{[m,1]}:S_{m}\to S_{m}$ is the algebra homomorphism determined by $\psi_{[m,1]}(a)=a^{q}-ax_{m}^{q-1}$ for $a$ homogeneous of degree one (see Section 2). Using Lemma 4.2, $x_{m}S_{m}=\ker(\psi_{[m,1]})$. Let $\overline{\psi_{[m,1]}}$ denote the algebra map from $A/x_{m}A$ to $S_{m}$ induced by $\psi_{[m,1]}$. If $\overline{\psi_{[m,1]}}$ is injective, then $A/x_{m}A$ is an integral domain, proving $x_{m}A$ is a prime ideal and $A$ is normal.
@@
Define $\overline{Q}:=\mathbb{F}_{q}[N(y_{m}),Y_{m-1},\ldots,Y_{1},X_{1},\ldots,X_{m-1}]$ (the subalgebra generated by $\{N(y_{m})\}\cup\mathcal{W}$). Let $\mathcal{B}_{H}$ denote the set of monomial factors of $(\xi_{1}\cdots\xi_{n-2})^{q-1}$. Using the $Q^{H}$-module structure of $A$, we see that $A/x_{m}A$ is a free $\overline{Q}$-module with a basis $\mathcal{B}_{H}$. By identifying $\overline{Q}$ with the submodule of $A/x_{m}A$ generated by $1$ we see that $A/x_{m}A$ is a complete intersection with relations $\xi_{i}^{q}=\psi_{[m,1]}(\xi_{i})$ for $i=1,\ldots,n-2$. From Lemma 4.2, for $f\in S_{m}$, $\psi_{[m,1]}\circ\psi_{[m,1]}(f)=\psi_{[m,1]}(f)^{q}$. Therefore $\psi_{[m,1]}(A)$ is a module over $\mathbb{F}_{q}[\psi_{[m,1]}(N(y_{m})),Y_{m-1}^{q},\dots,X_{m-1}^{q}]$ with module generators $\{\psi_{[m,1]}(\beta)\mid\beta\in\mathcal{B}_{H}\}$. To prove that $\overline{\psi_{[m,1]}}$ is injective, it is sufficient to show that $\overline{\psi_{[m,1]}}(A/x_{m}A)=\psi_{[m,1]}(A)$ is a free module over
$\psi_{[m,1]}(\overline{Q})=\mathbb{F}_{q}[\psi_{[m,1]}(N(y_{m})),Y_{m-1}^{q},\dots,X_{m-1}^{q}]$
with module generators $\{\psi_{[m,1]}(\beta)\mid\beta\in\mathcal{B}_{H}\}$. Note that for $\beta\in\mathcal{B}_{H}$, we have $\psi_{[m,1]}(\beta)\in\mathbb{F}_{q}[Y_{m-1}^{q},\dots,X_{m-1}^{q}]$. Therefore, it is sufficient to show that the module generated by $\{\psi_{[m,1]}(\beta)\mid\beta\in\mathcal{B}_{H}\}$ over $\mathbb{F}_{q}[Y_{m-1}^{q},\dots,X_{m-1}^{q}]$ is free. This is equivalent to showing the $\mathbb{F}_{q}[y_{m-1}^{q},\dots,x_{m-1}^{q}]$-module generated by $\{\xi_{1}^{e_{1}}\cdots\xi_{n-2}^{e_{n-2}}\mid e_{i}<q\}\subset S_{m-1}$ is a free module of rank $q^{n-2}$. This follows from the next lemma. ∎
@@
Suppose $G$ is a subgroup of $\text{GL}_{{n}}(\mathbb{F}_{q})$}} and the invariant field $\mathcal{F}(S_{m}^{G})$ is generated by $\{f_{1},\ldots,f_{n}\}\subset S_{m}^{G}$. Then the $\mathbb{F}_{q}[y_{m}^{q},\dots,y_{1}^{q},x_{1}^{q},\ldots,x_{m}^{q}]$-module generated by $\{f_{1}^{e_{1}}\cdots f_{n}^{e_{n}}\mid e_{i}<q\}\subset S_{m}$ is a free module of rank $q^{n}$.
@@
Since the extension $\mathbb{F}_{q}(f_{1},\ldots,f_{n})=\mathcal{F}(S_{m})^{G}\subset\mathcal{F}(S_{m})$ is a Galois extension, for any intermediate extension $\mathcal{F}(S_{m})^{G}\subset E\subset\mathcal{F}(S_{m})$, we have $E=\mathcal{F}(S_{m})^{L}$ for some subgroup $L\leq G$. Consider $E=\mathcal{F}(S_{m})^{G}[y_{m}^{q},...,x_{m}^{q}]$. Since the pointwise stabiliser of $\{y_{m}^{q},...,x_{m}^{q}\}$ in $\text{GL}_{{n}}(\mathbb{F}_{q})$}} is $\{1\}$, we have $\mathcal{F}(S_{m})^{G}[y_{m}^{q},...,x_{m}^{q}]=\mathcal{F}(S_{m})$. On the other hand, $\mathbb{F}_{q}(y_{m}^{q},...,x_{m}^{q})\subset\mathcal{F}(S_{m})$ is a field extension of degree $q^{n}$. Construct this extension by iteratively adjoining the $f_{i}$. Since $f_{i}$ is a root of $t^{q}-f_{i}^{q}\in\mathbb{F}_{q}(y_{m}^{q},...,x_{m}^{q})[t]$, at each step the extension has degree at most $q$. To complete the construction in $n$ steps, we need each step to be an extension of degree $q$. We can adjoin the $f_{i}$ in any order. Therefore $t^{q}-f_{i}^{q}$ is irreducible over $\mathbb{F}_{q}(y_{m}^{q},...,x_{m}^{q})[f_{j}\mid j\not=i]$. Hence $\{f_{1}^{e_{1}}\cdots f_{n}^{e_{n}}\mid e_{i}<q\}$ is a basis for $\mathcal{F}(S_{m})$ as a vector space over $\mathbb{F}_{q}(y_{m}^{q},...,x_{m}^{q})$. Thus $\{f_{1}^{e_{1}}\cdots f_{n}^{e_{n}}\mid e_{i}<q\}$ is linearly independent over $\mathbb{F}_{q}(y_{m}^{q},...,x_{m}^{q})$ and generates a free $\mathbb{F}_{q}[y_{m}^{q},...,x_{m}^{q}]$-module of rank $q^{n}$. ∎
[…]
Lemma 5.7.
If $f\in S^{H}$ is strongly compliant and $\deg_{y_{m}}(f)<q^{n-2}$, then $f\in Q^{H}[\xi_{1},\xi_{2},\ldots,\xi_{n-2}]$.
Proof.
The proof is by induction on $\ell=\deg_{y_{m}}(f)$. For $\ell=0$, $f=r_{0}\in Q^{H}$. Suppose the result is true for all strongly compliant invariants with $y_{m}$-degree $d$ and $0\leq d<\ell<q^{n-2}$.
[…]
We will prove by induction on $m$ that $N(y_{m})$ is $(y_{m},x_{m})$-compliant. As a consequence, $x_{m}N(y_{m})-\xi_{n-1}$ is strongly compliant. Lemma 5.3 then follows from Lemma 5.7. For $m=1$, we have $N(y_{1})=y_{1}$, which is $(y_{1},x_{1})$-compliant.
@@
Suppose that q1q-1 divides $k$, then q1q-1 divides $k||_{q}$.
@@
Write $k=\sum_{s}k_{s}q^{s}$ where $0\leq k_{s}\leq q-1$ for all $s$. Since $q^{s}\equiv 1\pmod{(q-1)}$ for all $s$ we have
[…]
@@
[…]
Given $f(y_{m},x_{1})\in T$, we define
$\mathcal{T}_{t}(f):=\prod_{a\in\mathbb{F}_{q}}f(y_{m}+at,x_{1}-ax_{m})\in T\ .$
@@
If $f$ is $(y_{m},x_{m})$-compliant then $\mathcal{T}_{t}(f)$ is $(y_{m},x_{m})$-compliant.